# 3D-GENERALIST: Vision-Language-Action Models for Crafting 3D Worlds

### v1

@@ Abstract
Despite large-scale pretraining endowing models with language and vision reasoning capabilities, improving their spatial reasoning capability remains challenging due to the lack of data grounded in the 3D world. While it is possible for humans to manually create immersive and interactive worlds through 3D graphics, as seen in applications such as VR, gaming, and robotics, this process remains highly labor-intensive. In this paper, we propose an scalable method for generating high-quality 3D environments that can serve as training data for foundation models. We recast 3D environment building as a sequential decision-making problem, employing Vision-Language-Models (VLMs) as policies that output actions to jointly craft a 3D environment’s layout, materials, lighting, and assets. Our proposed framework, 3D-Generalist, trains VLMs to generate more prompt-aligned 3D environments via self-improvement fine-tuning. We demonstrate the effectiveness of 3D-Generalist and the proposed training strategy in generating simulation-ready 3D environments. Furthermore, we demonstrate its quality and scalability in synthetic data generation by pretraining a vision foundation model on the generated data. After fine-tuning the pre-trained model on downstream tasks, we show that it surpasses models pre-trained on meticulously human-crafted synthetic data and approaches results achieved with real data orders of magnitude larger.
@@ 3.2 Scene-Level Policy
Let $S_{0}$ represent the 3D environment we obtained from Panoramic Environment Generation, $S_{t}$ be the 3D environment at iteration $t$, and let $\mathcal{P}$ denote the input text prompt. Our Vision-Language Model (VLM), $\pi_{\theta}$, parameterized by $\theta$, takes multi-view images $\mathcal{I}_{t}$ of $S_{t}$ (as shown in Figure 3) along with $\mathcal{P}$ as input and outputs an action code $a_{t}\sim\pi_{\theta}\bigl{(}a\mid\mathcal{I}_{t},\mathcal{P}\bigr{)}$. This action code is executed using exposed tools and function APIs—such as those for retrieving assets or materials based on natural language—which updates the 3D environment, expressed as $S_{t+1}=f\bigl{(}S_{t},a_{t}\bigr{)}$. Following the update, new images $\mathcal{I}_{t+1}$ are rendered from $S_{t+1}$ to serve as the next state representation. In Figure 4’s example, the action code, upon execution, adjusts the position of the bathtub, adds a bathroom vanity and two plants, and increases the scene’s lighting. We use GPT-4o [1] as the base model.
@@
Off-the-shelf VLMs often struggle to consistently generate action code that improves the resulting 3D environment. To address this, we propose a self-improvement fine-tuning strategy. Let $\pi^{(i)}_{\theta}$ denote the VLM policy at the start of the $i$-th fine-tuning round. Each round begins with generating a set of tasks (i.e., text prompts). For each task, the 3D environment is initialized using Panoramic Environment Generation, and $\pi^{(i)}_{\theta}$ iteratively updates the environment. At each state, multiple candidate actions are generated, and the top-scoring action sequences—those that yield the highest CLIP scores between the environment renderings and the input prompt—are retained. These top actions are then used to update the policy parameters via supervised fine-tuning, resulting in an improved policy $\pi^{(i+1)}_{\theta}$.
This process completes one self-improvement round, after which $\pi^{(i+1)}_{\theta}$ can generate new tasks and environments for further refinement in subsequent rounds. Additionally, we introduce an in-context library, a collection of action codes that significantly enhances CLIP alignment scores. Empirically, we find that leveraging the in-context library during data generation substantially improves performance by promoting diversity and increasing the likelihood of discovering effective action sequences. More details can be found in the supplementary material.
@@ 3.3 Asset-Level Policy
Our scene-level VLM policy, used in Scene-Level Policy, often omits smaller objects (e.g., books, plates, utensils) and focuses on larger and more defining environmental assets. In contrast, placing small assets requires a different design approach to ensure strong physical plausibility. To address this, we introduce Asset-Level Policy, which refines the environment by composing smaller assets with “receptacle objects” (e.g., shelves, tables, counters), closely mimicking the construction process of physically grounded environments. Additionally, Asset-Level Policy is equipped with physical plausibility verifiers applied prior to placement. The system is inherently extensible, supporting custom plug-in verifiers such as physics engines or logic-based constraints.
The process begins with GPT-4o determining whether a 3D asset qualifies as a “receptacle object” that can host smaller items. Once identified as a receptacle, we initiate an iterative process to place these smaller assets. Before the first round, we have a base object $\mathcal{O}$ and a text prompt $\mathcal{P}$ (“wooden office desk in a writer’s home”). Below, we describe the process at the $k$-th round:
@@ 4 Experiments
Given 3D-Generalist’s iterative nature and our goal of generating a large-scale dataset for rendering over 10 million images in experiment (d), we make design choices (see supplementary) to stay within a feasible time and compute budget.
@@ 4.1 Simulation-Ready 3D Environment Generation
Following existing works [34, 20, 61, 59], we evaluate simulation-ready 3D environments by measuring semantic and prompt alignment using CLIP scores and complementing this with a human study to address CLIP’s limitations.
Comparative Study
To answer (a), we compare with Holodeck [55], the state-of-the-art in sim-ready 3D environment generation, as other methods lack prompt-driven or simulation-ready outputs. 3D-Generalist outperforms Holodeck using the same evaluation protocol (CLIP score over 250 prompts across 50 room types). Ablations show that the in-context example library and self-improving fine-tuning are crucial to Scene-Level Policy’s effectiveness.
@@
Table 2: Percentage of times human raters preferred each method.
Metric Holodeck 3D-Generalist
Prompt Alignment 21% 79%
Layout Coherence 20% 80%
Overall Preference 18% 82%
Physics Semantics Overall Score
CF IB Pos. Rot. PSA
LayoutGPT 83.8 24.2 80.8 78.0 16.6
Holodeck 77.8 8.1 62.8 55.6 5.6
I-Design 76.8 34.3 68.3 62.8 18.0
LayoutVLM 81.8 94.9 77.5 73.2 58.8
3D-Generalist 100.0 100.0 78.2 79.1 68.8
@@
To answer (b), we assess our self-improving fine-tuning strategy using quantitative metrics (Table 1, 3) and qualitative analysis (Figure 6). In quantitative metrics, we evaluate 3D-Generalist on both CLIP scores and physical plausibility. The x-axis in Figure 6 represents ”rounds of action” within a generation sequence. While GPT-4o (baseline) shows some early refinement, our fine-tuned model demonstrates significantly stronger iterative improvement. Qualitative examples highlight 3D-Generalist’s Scene-Level Policy’s self-correction. In a human study (Table 2), five graduate students evaluated 100 room generations (two views each) on prompt alignment, layout coherence, and preference. With five graduate students conducting the evaluation, we find that 3D-Generalist is preferred over Holodeck in approximately 4 out of 5 generations across all three metrics.
@@
In our experiments, GPT-4o is responsible for determining whether a 3D asset qualifies as a receptacle object, while another VLM, Molmo-7B [12], provides the visual cues for pixel-level placement. As shown in Figure 5 5’s qualitative examples, Asset-Level Policy can place books on shelves, stack plates on tables, or fill open shelves with suitable items.
@@
We further discuss and show qualitative example of the importance of leveraging the semantic knowledge in the VLM (i.e., Molmo) to predict asset placement in the supplementary material.
@@
ProcTHOR [10] also has a method to place small objects on surfaces of “receptacle objects”. However, rather than being placed according to a natural language prompt, objects are placed based on the frequency that a given object type appears on the receptacle in the hand-modeled AI2thor and RoboTHOR datasets. Thus, we opt to compare with a baseline inspired by ARCHITECT’s [50] small object placement that uses an inpainting model. Please find the details of our implementation in the supplementary material. After placing the objects, all methods are evaluated with the same CLIP similarity score.
Across the 15 test cases, the base receptacle objects (without any other objects) have an average CLIP score of 0.264. The baseline method produces composed assets with an average CLIP score of $0.269\pm 0.001$, and our Asset-Level Policy produces an average score of $0.282\pm 0.001$ , demonstrating the ability to place assets on “receptacle assets” in a prompt-aligned and physically plausible way  [^1] During our evaluation process, an asset can only be placed if the resulting scene is physically plausible (i.e., no mesh collision). Qualitatively, we observe that the baseline inpainting-based method regularly generates many overlapping objects and cannot place an item behind another. Our method, however, has access to multiple views during the iterative process and can thus spread out the object placement more naturally. Refer to Figure 7 for qualitative comparisons.
@@ 4.2 Downstream Applications
To demonstrate 3D-Generalist’s capability in scaling up 3D data, we conduct large-scale pretraining of vision encoders on the generated synthetic data. We render millions of images from generated scenes using Omniverse to create a large-scale dataset and compare feature extractors trained on our synthetic data with those trained on HyperSim [37], an artist-crafted hyper-realistic dataset, along with a real-world dataset.
@@ 5 Conclusion
We introduce 3D-Generalist, a framework that formulates text-driven 3D world creation as a sequential decision-making problem. The framework can be extended to integrate with other generative methods, such as image-to-3D reconstruction [44, 52], image-to-PBR material generation [27] or 3D asset generation models [30], enabling a fully generative graphics pipeline. Beyond content creation, the use of explicit 3D representations also facilitates generative simulation—enabling the automatic generation of environments for training RL policies [43, 49]. Limitations of 3D-Generalist are further discussed in the supplementary material.
Despite the demonstrated efficacy of 3D-Generalist in scaling up the generation of data grounded in 3D environments, 3D-Generalist has its limitations in generating functional layouts. An important future direction is to address how we can better generate layouts of a 3D space and consider the functional and contextual use of objects in a space based on the given prompt.
Acknowledgments.
We particularly thank Sharon Lee, Rao Fu, and Jianhe Yuan for useful discussions, as well as Ritesh Chavan and Kevin Kuo for their support.
@@ 6.1 Scene-Level Policy
Here, we detail the types of descriptors defined in our Scene DSL. The Category descriptor, $C$, specifies the type of the scene element, where $C\in\{\text{floors},\text{walls},\text{ceilings},\text{objects}\}$. The Placement descriptor, $P=[(x,y,z),\theta_{z},s]$, defines the spatial attributes of the element, where $(x,y,z)\in\mathbb{R}^{3}$ specifies the position in 3D space, $\theta_{z}\in[0,2\pi]$ represents the rotation around the z-axis, and $s\in\mathbb{R}^{3}$ is the scale of the element. Placement can include a list of such tuples to allow for multiple placements of a particular scene element. The Material descriptor, $M=\text{desc}$, provides a natural language description ($\text{desc}\in\text{NL}$) of the element’s surface properties. The Lighting descriptor, $L=(t,i,c)$, specifies the type of the light, $t\in\{\text{point},\text{directional},\text{area}\}$, along with its intensity $i\in\mathbb{R}^{+}$ and color $c\in\mathbb{R}^{3}$, where $c=(r,g,b)$ and $r,g,b\in[0,1]$ represent the normalized RGB color values. Finally, the Metadata descriptor, $D$, captures additional category-specific attributes, details of which are provided in the supplementary materials. Below, we provide the implementation in Python:
@@ 6.2 Asset-Level Policy
6.3 Limitations & Additional Qualitative Results
A key limitation of 3D-Generalist is that it is fine-tuned to generate prompt-aligned 3D environments measure by CLIP similarity score, which is a metric not inherently sensitive to different spatial arrangements of the same set of assets. As a result, generated scenes often lack nuanced spatial reasoning, leading to suboptimal layout.Additional examples of 3D-Generalist’s Asset-Level Policy are in Figure 9.
@@ 7.3 Asset-Level Policy’s Evaluation
Baseline Implementation
@@
Figure 9: More qualitative results from 3D-Generalist’s Asset-Level Policy.

### v2

@@ Abstract
Creating 3D graphics content for immersive and interactive worlds remains labor-intensive, limiting our ability to create large-scale synthetic data that can serve as training data to foundation models. Recent methods have been proposed to alleviate this, but they often focus on one particular aspect (e.g., layout) and fail to improve the quality of the generation through scaling computational resources. In this work, we recast 3D environment generation as a sequential decision-making problem, employing Vision-Language-Models (VLMs) as policies that output actions to jointly craft a 3D environment’s layout, materials, lighting, and assets. Our proposed framework, 3D-Generalist, trains VLMs to generate more prompt-aligned 3D environments via self-improvement fine-tuning. We demonstrate the effectiveness of 3D-Generalist and the proposed training strategy in generating simulation-ready 3D environments. Furthermore, we demonstrate its quality and scalability in synthetic data generation by pretraining a vision foundation model on the generated data. After fine-tuning the pre-trained model on downstream tasks, we show that it surpasses models pre-trained on meticulously human-crafted synthetic data and approaches results achieved with real data orders of magnitude larger.
@@ 3.2 Scene-Level Policy
Let $S_{0}$ represent the 3D environment we obtained from Panoramic Environment Generation, $S_{t}$ be the 3D environment at iteration $t$, and let $\mathcal{P}$ denote the input text prompt. Our VLM, $\pi_{\theta}$, parameterized by $\theta$, takes multi-view images $\mathcal{I}_{t}$ of $S_{t}$ (as shown in Figure 3) along with $\mathcal{P}$ as input and outputs an action code $a_{t}\sim\pi_{\theta}\bigl{(}a\mid\mathcal{I}_{t},\mathcal{P}\bigr{)}$. This action code is executed using exposed tools and function APIs—such as those for retrieving assets or materials based on natural language—which updates the 3D environment, expressed as $S_{t+1}=f\bigl{(}S_{t},a_{t}\bigr{)}$. Following the update, new images $\mathcal{I}_{t+1}$ are rendered from $S_{t+1}$ to serve as the next state representation. In Figure 4’s example, the action code, upon execution, adjusts the position of the bathtub, adds a bathroom vanity and two plants, and increases the scene’s lighting.
@@
Off-the-shelf VLMs often struggle to consistently generate action code that improves the resulting 3D environment. To address this, we propose a self-improvement fine-tuning strategy. Let $\pi^{(i)}_{\theta}$ denote the VLM policy at the start of the $i$-th fine-tuning round. Each round begins with generating a set of tasks (i.e., text prompts). For each task, the 3D environment is initialized using Panoramic Environment Generation, and $\pi^{(i)}_{\theta}$ iteratively updates the environment. At each state, multiple candidate actions are generated, and the top-scoring action sequences—those that yield the highest CLIP scores between the environment renderings and the input prompt—are retained. These top actions are then used to update the policy parameters via supervised fine-tuning, resulting in an improved policy $\pi^{(i+1)}_{\theta}$. This process completes one self-improvement round, after which $\pi^{(i+1)}_{\theta}$ can generate new tasks and environments for further refinement in subsequent rounds. Additionally, we introduce an in-context library, a collection of action codes that significantly enhances CLIP alignment scores. Empirically, we find that leveraging the in-context library during data generation substantially improves performance by promoting diversity and increasing the likelihood of discovering effective action sequences. We use GPT-4o [1] as the base model and leveraged its publicly available fine-tuning API in our experiments. More details can be found in the supplementary material.
@@ 3.3 Asset-Level Policy
Our scene-level VLM policy, used in Scene-Level Policy, often omits smaller objects (e.g., books, plates, utensils) and focuses on larger and more defining environmental assets. In contrast, placing small assets requires a different design approach to ensure strong physical plausibility. To address this, we introduce Asset-Level Policy, which refines the environment by composing smaller assets with “receptacle objects” (e.g., shelves, tables, counters), closely mimicking the construction process of physically grounded environments. Asset-Level Policy is equipped with physical plausibility verifiers applied prior to placement. The system is inherently extensible, supporting custom plug-in verifiers such as physics engines or logic-based constraints. The process begins with GPT-4o determining whether a 3D asset qualifies as a “receptacle object” that can host smaller items. Once identified as a receptacle, we initiate an iterative process to place these smaller assets. Before the first round, we have a base object $\mathcal{O}$ and a text prompt $\mathcal{P}$ (“wooden office desk in a writer’s home”). Below, we describe the process at the $k$-th round:
@@ 4.1 Simulation-Ready 3D Environment Generation
Following existing works [34, 20, 61, 59, 42], we evaluate simulation-ready 3D environments by measuring physical plausibility and semantic coherence. Physical plausibility is measured using the Collision-Free Score (CF) and In-Boundary Score (IB). All assets are enforced to be placed, with remaining assets randomly placed if a method fails. Semantic coherence is assessed using Positional Coherency (Pos.) and Rotational Coherency (Rot.), measuring alignment with the input prompt. To evaluate semantic coherence across layouts without groundtruth, we use GPT-4o to score layouts based on top-down and side-view renderings and the language instructions.Physically-Grounded Semantic Alignment Score (PSA) is calculated simply the GPT-4o rating weighted by physical plausibility. Scores range from 0 to 100, with higher scores indicating better performance. Note that we don’t enable gravity in all experiments, and that a collision is considered to occur when the intersecting area between two meshes exceeds 0.1 m² in Blender.
Physics Semantics Overall Score
CF IB Pos. Rot. PSA
LayoutGPT 83.8 24.2 80.8 78.0 16.6
Holodeck 77.8 8.1 62.8 55.6 5.6
I-Design 76.8 34.3 68.3 62.8 18.0
LayoutVLM 81.8 94.9 77.5 73.2 58.8
3D-Generalist 99.0 98.0 78.2 79.1 67.9
@@
Comparative Study
To answer (a), we choose LayoutGPT [16], Holodeck [55], and LayoutVLM [42], the state-of-the-art in sim-ready 3D environment generation methods as baselines. While there are many other works on 3D scene generation like Physcene [54] and DiffuScene [46], they are either limited to a fixed set of object categories (i.e., not open-vocabulary methods) or rely on representations such as Gaussian or NeRF as opposed to resorting to reasoning models (i.e., LLMs, VLMs) with object-centric representation. 3D-Generalist outperforms all baselines using the same set of prompts used in Holodeck [55] (over 250 prompts across 50 room types).
@@
To answer (b), we assess our self-improving fine-tuning strategy using quantitative metrics and qualitative analysis (CLIP scores in Table 2 and Figure 6). The x-axis in Figure 6 represents ”rounds of action” within a generation sequence. While GPT-4o (baseline) shows some early refinement, our fine-tuned model demonstrates significantly stronger iterative improvement. Qualitative examples highlight Scene-Level Policy’s ability to self-correct.
Additional Ablation on Fine-tuning
@@
In our experiments, GPT-4o is responsible for determining whether a 3D asset qualifies as a receptacle object, while another VLM, Molmo-7B [12], provides the visual cues for pixel-level placement. As shown in Figure 5’s qualitative examples, Asset-Level Policy can place books on shelves, stack plates on tables, or fill open shelves with suitable items.
@@
Across the 15 test cases, the base receptacle objects (without any other objects) have an average CLIP score of 0.264. The baseline method produces composed assets with an average CLIP score of $0.269\pm 0.001$, and our Asset-Level Policy produces an average score of $0.282\pm 0.001$ , demonstrating the ability to place assets on “receptacle assets” in a prompt-aligned and physically plausible way  [^1] During our evaluation process, an asset can only be placed if the resulting scene is physically plausible (i.e., no mesh collision). Qualitatively, we observe that the baseline inpainting-based method regularly generates many overlapping objects and cannot place an item behind another. Our method, however, has access to multiple views during the iterative process and can thus spread out the object placement more naturally. Refer to Figure 7 and the supplementary material for qualitative comparisons that showcase of the importance of leveraging the semantic knowledge in the VLM (i.e., Molmo) to predict asset placement.
@@ 4.2 Downstream Applications
To demonstrate 3D-Generalist’s capability in scaling up 3D data, we conduct large-scale pretraining of vision encoders on the generated synthetic data. We render millions of images from generated scenes using Omniverse to create a large-scale dataset and compare feature extractors trained on our synthetic data with those trained on HyperSim [37], an artist-crafted dataset, along with a real-world dataset.
@@ 5 Conclusion
We introduce 3D-Generalist, a framework that formulates text-driven 3D world creation as a sequential decision-making problem. The framework can be integrated with image-to-3D reconstruction [44, 52], image-to-PBR material generation [27] or 3D asset generation models [30], enabling a fully generative graphics pipeline. Beyond content creation, the use of explicit 3D representations also facilitates generative simulation—enabling the automatic generation of environments for training RL policies [43, 49].
@@ 6.1 Scene-Level Policy
[…]
@@ 7.3 Asset-Level Policy’s Evaluation
ProcTHOR [10] also has a method to place small objects on surfaces of “receptacle objects”. However, rather than being placed according to a natural language prompt, objects are placed based on the frequency that a given object type appears on the receptacle in the hand-modeled AI2thor and RoboTHOR datasets. Thus, we opt to compare with an inpainting-based baseline mainly.
@@
7.4 GPU cost
Unlike existing monolithic pipelines, 3D-Generalist’s iterative design allows rendering and reasoning over complex scenes in a way that can scale with compute. All our experiments are conducted with A100s, with 1-3 minutes needed to generate the 3D scenes shown in Figure 1.